# Composing Codensity Bisimulations

### v1

@@ 8.1 Compositionality of Invariants
Let $((\Omega,\mathbf{\Omega}):Id_{A}\to(p:\mathbb{E}\to\mathbb{B}),F,\tau)$ be codensity bisimulation data and $c\colon X\to F(X)$ be a $F$-coalgebra. The codensity game $\mathcal{G}_{c}$ is the safety game $(Q_{\mathbf{D}},Q_{\mathbf{S}},E)$ played by two players $\mathbf{D},\mathbf{S}$ where $Q_{\mathbf{D}}\coloneqq\{(a,k)\mid a\in A,k\in\mathbb{B}(X,\Omega(a))\}$, $Q_{\mathbf{S}}\coloneqq\mathrm{Obj}(\mathbb{E}_{X})$, and $E\coloneqq{\to_{\mathbf{S}}}\cup{\to_{\mathbf{D}}}$ is given by
@@
    A set $\mathcal{V}\subseteq\mathrm{Obj}(\mathbb{E}_{X})$ is an invariant for $\mathbf{D}$ in the codensity game $\mathcal{G}_{c}$ if and only if $\bigsqcup_{P\in\mathcal{V}}P$ is a $[F,\tau]^{\mathbf{\Omega}}$-bisimulation on $c$.
@@
    A position $P\in\mathbb{E}_{X}$ is winning in $\mathcal{G}_{c}$ if and only if $P$ is a witness of the codensity bisimilarity on $c$.
@@
Let $((\Omega,\mathbf{\Omega}):Id_{A}\to(p:\mathbb{E}\to\mathbb{B}),F,\tau)$ be a codensity bisimulation data and $c_{i}\colon X_{i}\to F(X_{i})$ ($i\in N$) be $F$-coalgebras. The composite codensity game $\mathcal{G}_{\overrightarrow{{c_{i}}}}^{T,\sigma}$ is the safety game $(Q_{\mathbf{D}},Q_{\mathbf{S}},E)$ by two players $\mathbf{D},\mathbf{S}$ where $Q_{\mathbf{D}}\coloneqq\{(a,\overrightarrow{{k_{i}}})\mid a\in A,k_{i}\in\mathbb{C}(X_{i},\Omega(a))\}$, $Q_{\mathbf{S}}\coloneqq\prod_{i}\mathrm{Obj}(\mathbb{E}_{X_{i}})$, and $E\coloneqq\to_{\mathbf{S}}\cup\to_{\mathbf{D}}$ is given by
@@ 8.2 Preservation of Bisimilarities
Under the sufficient condition in Theorem 6.2, invariants of composite codensity games $\mathcal{G}_{\overrightarrow{{c_{i}}}}^{T,\sigma}$ also characterizes codensity bisimilarities on composite coalgebras ${T}_{\lambda}(\overrightarrow{{c_{i}}})$. We introduce a join-closure operation for invariants: for a set $\mathcal{V}\subseteq\prod_{i\in N}\mathrm{Obj}(\mathbb{E}_{X_{i}})$, we write $\overline{\mathcal{V}}$ for $\mathcal{V}\cup\Big{\{}\left({\bigsqcup_{\overrightarrow{{P_{i}}}\in\mathcal{V}}P_{i}}\right)_{i\in N}\Big{\}}$. We remark that if a set $\mathcal{V}$ is an invariant of a codensity game $\mathcal{G}_{c}$ for $\mathbf{D}$ then so is $\overline{\mathcal{V}}$.
@@
    $\bigsqcup_{\overrightarrow{{P_{i}}}\in\mathcal{V}}[T,\sigma]^{\mathbf{\Omega}}(\overrightarrow{{P_{i}}})$ is a $[F,\tau]^{\mathbf{\Omega}}$-bisimulation on ${T}_{\lambda}(\overrightarrow{{c_{i}}})$.
@@
    $\bigsqcup_{\overrightarrow{{P_{i}}}\in\overline{\mathcal{V}}}[T,\sigma]^{\mathbf{\Omega}}(\overrightarrow{{P_{i}}})$ is a $[F,\tau]^{\mathbf{\Omega}}$-bisimulation on ${T}_{\lambda}(\overrightarrow{{c_{i}}})$. ∎
@@
Assume that the sufficient condition in Theorem 6.2 holds. If $\overrightarrow{{P_{i}}}$ is a winning position for $\mathbf{D}$ with an invariant $\overline{\mathcal{V}}$ then $[T,\sigma]^{\mathbf{\Omega}}(\overrightarrow{{P_{i}}})$ is a witness on ${T}_{\lambda}(\overrightarrow{{c_{i}}})$. ∎
Finally, through a composite codensity game, we prove the compositionality of witnesses (and equivalently the preservation of bisimilarities). If each $P_{i}$ is a witness on the codensity bisimilarity on $c_{i}$, $P_{i}$ is a winning position for $\mathbf{D}$ in $\mathcal{G}_{c_{i}}$ by Proposition 8.6. Then Proposition 8.4 ensures the existence of invariants $\overline{\mathcal{V}}_{i}$ of $\mathcal{G}_{c_{i}}$ such that $P_{i}\in\overline{\mathcal{V}}_{i}$. Proposition 8.8 and $\prod_{i\in N}\overline{\mathcal{V}}_{i}=\overline{\prod_{i\in N}\mathcal{V}_{i}}$ imply that $\overrightarrow{{P_{i}}}$ is a winning position with the invariant $\overline{\prod_{i\in N}\mathcal{V}_{i}}$. Therefore, Corollary 8.10 concludes that $[T,\sigma]^{\mathbf{\Omega}}(\overrightarrow{{P_{i}}})$ is a witness of the codensity bisimilarity on ${T}_{\lambda}(\overrightarrow{{c_{i}}})$.
@@ 9 Conclusion
For future work, we will continue to explore and expand the scope of our approach, particularly considering different structure functors or different notions of compatibility as they appear in the study of up-to techniques.
Recently, higher-order abstract GSOS frameworks [14, 33, 15] are actively studied: they extend behavior endofunctors to bifunctors and use certain dinatural transformations for distributive laws. Seeking a higher-order extension of our framework, that is, studying a liftability of distributive laws along codensity liftings in a higher-order setting is an interesting future direction.
@@ Acknowledgement
MK, KW and SK were supported by ERATO HASUO Metamathematics for Systems Design Project (No. JPMJER1603), JST. MK was supported by JSPS DC KAKENHI Grant (No. 22J21742). KW was supported by the JST grants No. JPMJFS2136 and JPMJAX23CU. JR was supported by the NWO grant No. OCENW.M20.053.
@@ A.1 Omitted Proofs for Section 5
In Example. 5.2, the ${\mathbf{CLat}_{\sqcap}}$-fibrations with truth values $(2,(2,\mathrm{Eq_{2}})):Id_{1}\to U_{\mathbf{EqRel}}$ and $(2,(2,\{\emptyset,\{\mathrm{true}\},2\})):Id_{1}\to U_{\mathbf{Top}}$ satisfy $R^{p,\mathbf{\Omega}}L^{p,\mathbf{\Omega}}=\mathrm{id}$.
@@
  • ($(2,(2,\mathrm{Eq_{2}})):Id_{1}\to U_{\mathbf{EqRel}}$): $\bigsqcap_{k\colon\mathbf{\Omega}\to\mathbf{\Omega}}k^{*}\mathbf{\Omega}=\{(x,y)\mid\forall k\colon\mathbf{\Omega}\to\mathbf{\Omega}.~{}kx=ky\}=\mathbf{\Omega}$ holds. The last equality is because $(x,y)$ in l.h.s. satisfies $x=y$ by letting $k\coloneqq\mathrm{id}$, and $(x,y)$ in r.h.s. satisfies $kx=ky$ for each $k\colon\mathbf{\Omega}\to\mathbf{\Omega}$.
  • ($(2,(2,\{\emptyset,\{\mathrm{true}\},2\})):Id_{1}\to U_{\mathbf{Top}}$): $\bigsqcap_{k\colon\mathbf{\Omega}\to\mathbf{\Omega}}k^{*}\mathbf{\Omega}$ is the topology generated by $\bigcup_{k\colon\mathbf{\Omega}\to\mathbf{\Omega}}\{k^{-1}A\mid A\in\mathbf{\Omega}\}$. $\mathbf{\Omega}$ is included in the topology by $k\coloneqq\mathrm{id}\colon\mathbf{\Omega}\to\mathbf{\Omega}$, and the topology is included in $\mathbf{\Omega}$ because $k^{-1}A\in\mathbf{\Omega}$ for each $k\colon\mathbf{\Omega}\to\mathbf{\Omega}$ and $A\in\mathbf{\Omega}$.
@@ A.3.1 Bisimilarity Pseudometric
In the setting of §7.1, the tuple $((\Omega,\mathbf{\Omega}):Id_{A}\to U_{\mathbf{PMet}},(\times,\lambda^{\mathrm{da}}),\sigma_{\land})$ satisfies the conditions in Theorem 6.2.
@@ A.3.2 Similarity Pseudometric
    Fix arbitrary $d_{1},d_{2}\in\mathbf{LMet}_{X}$ and $(x,y),(x^{\prime},y^{\prime})\in X\times Y$. Let us prove [×$[\times,\sigma_{\land}]^{d_{\mathbb{I}}^{\mathrm{as}}}(d_{1},d_{2})((x,y),(x^{\prime},y^{\prime}))=\max(d_{1}(x,x^{\prime}),d_{2}(y,y^{\prime}))$. ()($\geq$)( ≥ ): Letting $k_{1}(t)\coloneqq 1-d_{1}(t,x^{\prime})$ and $k_{2}(t)\coloneqq 1-d_{2}(t,y^{\prime})$, $k_{1}\colon d_{1}\to d_{\mathbb{I}}^{\mathrm{as}}$ and $k_{2}\colon d_{2}\to d_{\mathbb{I}}^{\mathrm{as}}$. Therefore we have
    $(lhs)$ $\geq d_{\mathbb{I}}^{\mathrm{as}}(\min(k_{1}x,k_{2}y),\min(k_{1}x^{\prime},k_{2}y^{\prime}))$
    $=\max(d_{1}(x,x^{\prime}),d_{2}(y,y^{\prime}))=(rhs).$
    ()($\leq$)( ≤ ): For each $k_{1}\colon d_{1}\to d_{\mathbb{I}}^{\mathrm{as}}$ and $k_{2}\colon d_{2}\to d_{\mathbb{I}}^{\mathrm{as}}$,
    $d_{\mathbb{I}}^{\mathrm{as}}(\min(k_{1}x,k_{2}y),\min(k_{1}x^{\prime},k_{2}y^{\prime}))$
    $\leq\min(k_{1}x^{\prime},k_{2}y^{\prime})-\min(k_{1}x,k_{2}y)$
    $\leq\max(k_{1}x^{\prime}-k_{1}x,k_{2}y^{\prime}-k_{2}y)$
    $\leq\max(d_{1}(x,x^{\prime}),d_{2}(y,y^{\prime}))$
    by Lemma A.5. Therefore we have (lhs) $\leq$ (rhs).
@@
In the setting of §7.2, the tuple $((\Omega,\mathbf{\Omega}):Id_{A}\to(p:\mathbf{LMet}\to\mathbf{Set}),(\times,\lambda^{\mathrm{da}}),\sigma_{\land})$ satisfies the conditions in Theorem 6.2.

### v2

@@ 8.1 Compositionality of Invariants
Let $((\Omega,\mathbf{\Omega})\colon Id_{A}\to(p\colon\mathbb{E}\to\mathbb{B}),F,\tau)$ be codensity bisimulation data and $c\colon X\to F(X)$ be an $F$-coalgebra. The codensity game $\mathcal{G}_{c}$ is the safety game $(Q_{\mathbf{D}},Q_{\mathbf{S}},E)$ played by two players $\mathbf{D},\mathbf{S}$ where $Q_{\mathbf{D}}\coloneqq\{(a,k)\mid a\in A,k\in\mathbb{B}(X,\Omega(a))\}$, $Q_{\mathbf{S}}\coloneqq\mathrm{Obj}(\mathbb{E}_{X})$, and $E\coloneqq{\to_{\mathbf{S}}}\cup{\to_{\mathbf{D}}}$ is given by
@@
    A set $\mathcal{V}\subseteq\mathrm{Obj}(\mathbb{E}_{X})$ is an invariant for $\mathbf{D}$ in the codensity game $\mathcal{G}_{c}$ if and only if $\bigsqcup_{P\in\mathcal{V}}P$ is an $[F,\tau]^{\mathbf{\Omega}}$-bisimulation on $c$.
@@
    A position $P\in\mathbb{E}_{X}$ is winning in $\mathcal{G}_{c}$ if and only if $P$ is a witness of codensity bisimilarity on $c$, that is, $P\sqsubseteq\nu\bigl{(}c^{*}\circ[F,\tau]^{\mathbf{\Omega}}\bigr{)}$. ∎
@@
Let $((\Omega,\mathbf{\Omega})\colon Id_{A}\to(p\colon\mathbb{E}\to\mathbb{B}),F,\tau)$ be a codensity bisimulation data and $c_{i}\colon X_{i}\to F(X_{i})$ ($i\in N$) be $F$-coalgebras. The composite codensity game $\mathcal{G}_{\overrightarrow{{c_{i}}}}^{T,\sigma}$ is the safety game $(Q_{\mathbf{D}},Q_{\mathbf{S}},E)$ by two players $\mathbf{D},\mathbf{S}$ where $Q_{\mathbf{D}}\coloneqq\{(a,\overrightarrow{{k_{i}}})\mid a\in A,k_{i}\in\mathbb{C}(X_{i},\Omega(a))\}$, $Q_{\mathbf{S}}\coloneqq\prod_{i}\mathrm{Obj}(\mathbb{E}_{X_{i}})$, and $E\coloneqq\to_{\mathbf{S}}\cup\to_{\mathbf{D}}$ is given by
@@
Example 8.8.
Consider language equivalence of deterministic automata as presented in §6.1.2. For an ${F_{\mathrm{da}}}$-coalgebra $c\colon X\to{F_{\mathrm{da}}}X$, the codensity game $\mathcal{G}_{c}$ is $((\Sigma\uplus\{\epsilon\})\times\mathbf{Set}(X,2),\mathrm{Obj}(\mathrm{EqRel}_{X}),\to_{\mathbf{S}}\cup\to_{\mathbf{D}})$ with
$P\to_{\mathbf{S}}(\epsilon,k)\text{ if }\exists(x,y)\in P.~{}\pi_{1}c(x)\neq\pi_{1}c(y),$
$P\to_{\mathbf{S}}(a\in\Sigma,k)\text{ if }\exists(x,y)\in P.~{}k(\pi_{2}c(x))\neq k(\pi_{2}c(y)),$
$(a,k)\to_{\mathbf{D}}P^{\prime}\text{ if }\exists(x,y)\in P^{\prime}.~{}k(x)\neq k(y),$
and for two ${F_{\mathrm{da}}}$-coalgebras $\{c_{i}\colon X_{i}\to{F_{\mathrm{da}}}X_{i}\}_{i\in 2}$, the composite codensity game $\mathcal{G}_{\overrightarrow{{c_{i}}}}^{\times,\sigma_{\land}}$ is $((\Sigma\uplus\{\epsilon\})\times\mathbf{Set}(X_{1},2)\times\mathbf{Set}(X_{2},2),\mathrm{Obj}(\mathrm{EqRel}_{X_{1}})\times\mathrm{Obj}(\mathrm{EqRel}_{X_{2}}),\to_{\mathbf{S}}\cup\to_{\mathbf{D}})$ with
$(P_{1},P_{2})\to_{\mathbf{S}}(\epsilon,k_{1},k_{2})$
[…]
$(P_{1},P_{2})\to_{\mathbf{S}}(a\in\Sigma,k_{1},k_{2})$
[…]
$(a,k_{1},k_{2})\to_{\mathbf{D}}(P^{\prime}_{1},P^{\prime}_{2})\text{ if }\exists i\in 2.~{}\exists(x,y)\in P^{\prime}_{i}.~{}k(x_{i})\neq k(y_{i}).$
In this composite codensity game, positions are composed by pairing and conditions are composed by the modality $\sigma_{\land}$.
@@ 8.2 Preservation of Bisimilarities
Under the sufficient condition in Theorem 6.2, as discussed in §6, we naturally have the following natural transformation above $\lambda$.
$R\circ\mathrm{Sp}^{A}(T,\sigma)\circ\mathrm{Sp}^{A}(F^{N},\tau^{N})\circ L_{N}\hfill$
[…]
It connects invariants in composite codensity games $\mathcal{G}_{\overrightarrow{{c_{i}}}}^{T,\sigma}$, which utilizes combination of modalities $\sigma\circ T(\vec{\tau})$, and $[F,\tau]^{\mathbf{\Omega}}$-bisimulations composed by $[T,\sigma]^{\mathbf{\Omega}}$. We introduce a join-closure operation for invariants: for a set $\mathcal{V}\subseteq\prod_{i\in N}\mathrm{Obj}(\mathbb{E}_{X_{i}})$, we write $\overline{\mathcal{V}}$ for $\mathcal{V}\cup\Big{\{}\left({\bigsqcup_{\overrightarrow{{P_{i}}}\in\mathcal{V}}P_{i}}\right)_{i\in N}\Big{\}}$. We remark that if a set $\mathcal{V}$ is an invariant of a codensity game $\mathcal{G}_{c}$ for $\mathbf{D}$ then so is $\overline{\mathcal{V}}$.
@@
    $\bigsqcup_{\overrightarrow{{P_{i}}}\in\mathcal{V}}[T,\sigma]^{\mathbf{\Omega}}(\overrightarrow{{P_{i}}})$ is an $[F,\tau]^{\mathbf{\Omega}}$-bisimulation on ${T}_{\lambda}(\overrightarrow{{c_{i}}})$.
@@
    $\bigsqcup_{\overrightarrow{{P_{i}}}\in\overline{\mathcal{V}}}[T,\sigma]^{\mathbf{\Omega}}(\overrightarrow{{P_{i}}})$ is an $[F,\tau]^{\mathbf{\Omega}}$-bisimulation on ${T}_{\lambda}(\overrightarrow{{c_{i}}})$. ∎
@@
Assume that the sufficient condition in Theorem 6.2 holds. If $\overrightarrow{{P_{i}}}$ is a winning position for $\mathbf{D}$ with an invariant $\overline{\mathcal{V}}$ then $[T,\sigma]^{\mathbf{\Omega}}(\overrightarrow{{P_{i}}})$ is a witness on ${T}_{\lambda}(\overrightarrow{{c_{i}}})$, that is, $[T,\sigma]^{\mathbf{\Omega}}(\overrightarrow{{P_{i}}})\sqsubseteq\nu\Bigl{(}{\big{(}{T}_{\lambda}(\overrightarrow{{c_{i}}})\big{)}}^{*}\circ[F,\tau]^{\mathbf{\Omega}}\Bigr{)}$. ∎
Finally, through a composite codensity game, we prove the preservation of bisimilarities (the inequalities shown in Corollary 3.10). In the proof, we instead prove the following equivalent statement:
Proposition 8.11.
Assume that the sufficient condition in Theorem 6.2 holds. If $P_{i}$ is a witness on $c_{i}$, that is, $P_{i}\sqsubseteq\nu\Bigl{(}{c_{i}}^{*}\circ[F,\tau]^{\mathbf{\Omega}}\Bigr{)}$ for each $i$, then $[T,\sigma]^{\mathbf{\Omega}}(\overrightarrow{{P_{i}}})$ is a witness on $T_{\lambda}(c_{1},\cdots,c_{N})$, that is, $[T,\sigma]^{\mathbf{\Omega}}(\overrightarrow{{P_{i}}})\sqsubseteq\nu(T_{\lambda}(c_{1},\cdots,c_{N})^{*}\circ[F,\tau]^{\mathbf{\Omega}})$.
We outline the alternative proof of the proposition by composite codensity games. If each $P_{i}$ is a witness on codensity bisimilarity on $c_{i}$, $P_{i}$ is a winning position for $\mathbf{D}$ in $\mathcal{G}_{c_{i}}$ by Proposition 8.5. Then Proposition 8.3 ensures the existence of invariants $\overline{\mathcal{V}}_{i}$ of $\mathcal{G}_{c_{i}}$ such that $P_{i}\in\overline{\mathcal{V}}_{i}$. Proposition 8.7 and $\prod_{i\in N}\overline{\mathcal{V}}_{i}=\overline{\prod_{i\in N}\mathcal{V}_{i}}$ imply that $\overrightarrow{{P_{i}}}$ is a winning position with the invariant $\overline{\prod_{i\in N}\mathcal{V}_{i}}$. Therefore, by Corollary 8.10 we conclude that $[T,\sigma]^{\mathbf{\Omega}}(\overrightarrow{{P_{i}}})$ is a witness of codensity bisimilarity on ${T}_{\lambda}(\overrightarrow{{c_{i}}})$.
@@ 9 Conclusion
For future work, we will continue to explore and expand the scope of our approach, particularly considering different structure functors.
Additionally, we intend to streamline the current categorical notation using the syntax of categorical logic.
Recently, higher-order abstract GSOS frameworks [14, 33, 15] are actively studied: they extend behavior endofunctors to bifunctors and use certain dinatural transformations for distributive laws. Seeking a higher-order extension of our framework, that is, studying a liftability of distributive laws along codensity liftings in a higher-order setting is an interesting future direction. More generally, in the current paper we focused on compositionality for individual operations. Accomodating languages and calculi in our approach as defined, for instance, by (abstract) GSOS specifications, remains future work.
@@ Acknowledgement
The authors are gradeful to Ichiro Hasuo, and Noboru Isobe for helpful discussions, and to the anonymous reviewers for helpful feedback. MK, KW and SK were supported by ERATO HASUO Metamathematics for Systems Design Project (No. JPMJER1603), JST. MK was supported by JSPS DC KAKENHI Grant (No. 22J21742). KW was supported by the JST grants No. JPMJFS2136 and JPMJAX23CU. JR was supported by the NWO grant No. OCENW.M20.053.
@@ A.1 Omitted Proofs for Section 5
In Example. 5.2, the ${\mathbf{CLat}_{\sqcap}}$-fibrations with truth values $(2,(2,\mathrm{Eq_{2}}))\colon Id_{1}\to U_{\mathbf{EqRel}}$ and $(2,(2,\{\emptyset,\{\mathrm{true}\},2\}))\colon Id_{1}\to U_{\mathbf{Top}}$ satisfy $R^{p,\mathbf{\Omega}}L^{p,\mathbf{\Omega}}=\mathrm{id}$.
@@
  • ($(2,(2,\mathrm{Eq_{2}}))\colon Id_{1}\to U_{\mathbf{EqRel}}$): $\bigsqcap_{k\colon\mathbf{\Omega}\to\mathbf{\Omega}}k^{*}\mathbf{\Omega}=\{(x,y)\mid\forall k\colon\mathbf{\Omega}\to\mathbf{\Omega}.~{}kx=ky\}=\mathbf{\Omega}$ holds. The last equality is because $(x,y)$ in l.h.s. satisfies $x=y$ by letting $k\coloneqq\mathrm{id}$, and $(x,y)$ in r.h.s. satisfies $kx=ky$ for each $k\colon\mathbf{\Omega}\to\mathbf{\Omega}$.
  • ($(2,(2,\{\emptyset,\{\mathrm{true}\},2\}))\colon Id_{1}\to U_{\mathbf{Top}}$): $\bigsqcap_{k\colon\mathbf{\Omega}\to\mathbf{\Omega}}k^{*}\mathbf{\Omega}$ is the topology generated by $\bigcup_{k\colon\mathbf{\Omega}\to\mathbf{\Omega}}\{k^{-1}A\mid A\in\mathbf{\Omega}\}$. $\mathbf{\Omega}$ is included in the topology by $k\coloneqq\mathrm{id}\colon\mathbf{\Omega}\to\mathbf{\Omega}$, and the topology is included in $\mathbf{\Omega}$ because $k^{-1}A\in\mathbf{\Omega}$ for each $k\colon\mathbf{\Omega}\to\mathbf{\Omega}$ and $A\in\mathbf{\Omega}$.
@@ A.3.1 Bisimilarity Pseudometric
In the setting of §7.1, the tuple $((\Omega,\mathbf{\Omega})\colon Id_{A}\to U_{\mathbf{PMet}},(\times,\lambda^{\mathrm{da}}),\sigma_{\land})$ satisfies the conditions in Theorem 6.2.
@@ A.3.2 Similarity Pseudometric
    Fix arbitrary $d_{1},d_{2}\in\mathbf{LMet}_{X}$ and $(x,y),(x^{\prime},y^{\prime})\in X\times Y$. Let us prove [×$[\times,\sigma_{\land}]^{d_{\mathbb{I}}^{\mathrm{as}}}(d_{1},d_{2})((x,y),(x^{\prime},y^{\prime}))=\max(d_{1}(x,x^{\prime}),d_{2}(y,y^{\prime}))$. ()($\geq$)( ≥ ): Letting $k_{1}(t)\coloneqq 1-d_{1}(t,x^{\prime})$ and $k_{2}(t)\coloneqq 1-d_{2}(t,y^{\prime})$, $k_{1}\colon d_{1}\to d_{\mathbb{I}}^{\mathrm{as}}$ and $k_{2}\colon d_{2}\to d_{\mathbb{I}}^{\mathrm{as}}$. Therefore we have (lhs) $\geq d_{\mathbb{I}}^{\mathrm{as}}(\min(k_{1}x,k_{2}y),\min(k_{1}x^{\prime},k_{2}y^{\prime}))=\max(d_{1}(x,x^{\prime}),d_{2}(y,y^{\prime}))=$ (rhs). ()($\leq$)( ≤ ): For each $k_{1}\colon d_{1}\to d_{\mathbb{I}}^{\mathrm{as}}$ and $k_{2}\colon d_{2}\to d_{\mathbb{I}}^{\mathrm{as}}$, $d_{\mathbb{I}}^{\mathrm{as}}(\min(k_{1}x,k_{2}y),\min(k_{1}x^{\prime},k_{2}y^{\prime}))\leq\min(k_{1}x^{\prime},k_{2}y^{\prime})-\min(k_{1}x,k_{2}y)\leq\max(k_{1}x^{\prime}-k_{1}x,k_{2}y^{\prime}-k_{2}y)\leq\max(d_{1}(x,x^{\prime}),d_{2}(y,y^{\prime}))$ by Lemma A.5. Therefore we have (lhs) $\leq$ (rhs).
@@
In the setting of §7.2, the tuple $((\Omega,\mathbf{\Omega})\colon Id_{A}\to(p\colon\mathbf{LMet}\to\mathbf{Set}),(\times,\lambda^{\mathrm{da}}),\sigma_{\land})$ satisfies the conditions in Theorem 6.2.